OpenStudy (anonymous):

Find the complex zeros of the polynomial function. Write f in factored form. a.) F(x)= x^3-5x^2+11x-15 b.) f(x)=x^4+26x^2+25 c.) F(x) = 3x^4-10x^3-12x^2 +122x-39

OpenStudy (anonymous):

too hard just cheat

OpenStudy (anonymous):

for the first one, try \(f(1)\)

OpenStudy (anonymous):

\[1-5+11-15\] no that is not zero try \(f(-1)\)

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

no, ok try \(f(3)\)

OpenStudy (anonymous):

you get -18

OpenStudy (anonymous):

no you get 0

OpenStudy (anonymous):

oh yeah sorry i multiplied wrong

OpenStudy (anonymous):

\[f(x)=(x-3) (x^2-2 x+5)\]

OpenStudy (anonymous):

find the the other two zeros by setting \[x^2-2x+5=0\] and completing the square

OpenStudy (anonymous):

so the complex zeros are 3, 1-2i, and 1+2i

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well 3 is a real root, the other two are complex

OpenStudy (anonymous):

\[x^4+26x^2+25\] \[(x^2+1)(x^2+25)=0\] solve that one in a heartbeat

OpenStudy (anonymous):

it was wrong the right answer was ... (x-3)(x-1-2i)(x-1+2i)

OpenStudy (anonymous):

oh sorry it changed the question cause i got it wrong

OpenStudy (anonymous):

no it was not wrong they wanted you to write the polynomial in factored form, not just find the zeros

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

did you get the second one?

OpenStudy (anonymous):

yeah i got x=-i,i,-5i,5i

OpenStudy (anonymous):

ok good`

OpenStudy (anonymous):

so now how do I write that in factored form ?

OpenStudy (anonymous):

\[(x+i)(x-i)(x+5i)(x-5i)\]

OpenStudy (anonymous):

how did you get the first step?

OpenStudy (anonymous):

cause i'm trying to do the third question and i don't know how you approached it

OpenStudy (anonymous):

\[c.) F(x) = 3x^4-10x^3-12x^2 +122x-39\] is different because the second one has only \(x^4\) and \(x^2\)

OpenStudy (anonymous):

you just have to check what gets zero, it is a pain in the neck i would cheat http://www.wolframalpha.com/input/?i=3x^4-10x^3-12x^2+%2B122x-39

OpenStudy (anonymous):

so would it just be (3-2i)(3+2i)